An invoice clerk receives a bill $37.50 for ten blank ledger books and three special filing trays .on phoning the supplier ( after discovering the order had been written incorrectly), the clerk agrees to return three of the ledger books in return for the supplier sending an extra filing tray . Given that there willl be extra $2.50 to pay and that there is 10% discount for orders of 10 ledger books or more ;
a. Derive a linear equation in x and y to represent the components of the original revised invoices, where x is the price of a single ledger book and y is the price of a filling tray.
b. Solve the two equation simultaneously for x and y

Answer by ankor@dixie-net.com(22740)

An invoice clerk receives a bill $37.50 for ten blank ledger books and three special filing trays.
On phoning the supplier ( after discovering the order had been written incorrectly), the clerk agrees to return three of the ledger books in return for the supplier sending an extra filing tray.
Given that there will be extra $2.50 to pay and that there is 10% discount for orders of 10 ledger books or more;
:
a. Derive a linear equation in x and y to represent the components of the original & revised invoices, where x is the price of a single ledger book and y is the price of a filling tray.
The original bill had a 10% discount on 10 ledgers, therefore the paid for 9
9x + 3y = 37.50
The 2nd bill was $2.50 more. 3 ledgers were returned, 1 tray added, therefore:
7x + 4y = 40
b. Solve the two equation simultaneously for x and y
9x + 3y = 37.50
7x + 4y = 40.00
multiply the 1st equation by 4, the 2nd equation by 3
36x + 12y = 150
21x + 12y = 120
-----------------subtraction eliminates y, find x
15x + 0 = 30
x = 30/15
x = $2.00 for ledgers
Use the 2nd original equation to find y
7(2) + 4y = 40
4y = 40 - 14
y = 26/4
y = $6.50 for trays
:
:
Check solutions in the 1st original equation
9(2) + 3(6.50) =
18 + 19.50 = 37.50